OpenStudy (stuck-help):

METAL and FAN( if explained enough and is the correct answer) simplify the expression. state any excluded values4m-8/4-2m

OpenStudy (phi):

is this \[ \frac{4m-8}{4-2m} \]

OpenStudy (stuck-help):

yes

OpenStudy (phi):

do you know how to factor 4 from (4m-8) ?

OpenStudy (stuck-help):

no i dont know how to factor

OpenStudy (phi):

do you know about "distribute" as in 2( 3+4)= 2*3 + 2*4 ?

OpenStudy (stuck-help):

14

OpenStudy (phi):

yes both sides are 14. The important part is the "rule" if you have 2(3+4) you can think of that as a "package" of (3+4) and you have 2 packages, when you write 2(3+4) so you must have 2 of everything inside the package: 2*3 + 2*4

OpenStudy (stuck-help):

where did you get the 2

OpenStudy (phi):

It's an example. Your life will be easier if you learn this idea. another example, with letters 4*(m-2) that is 4 packages of (m-2), so we should expect 4 of each thing inside the package 4(m-2) is the same as 4*m + 4*-2

OpenStudy (stuck-help):

so 2(3+4) would be my answer

OpenStudy (stuck-help):

what about the 3 also i dont understand this

OpenStudy (phi):

no. But notice 4(m-2) = 4*m + 4*-2 do you see how I did that ?

OpenStudy (stuck-help):

yes i understand disturbuting but the answer to 4m-8/4-2m

OpenStudy (phi):

factoring is "undoing" distributing notice 4(m-2) = 4*m + 4*-2 we write 4*m as 4m (leave out the multiply sign) and 4*-2 is -8 in other words 4(m-2) = 4m + -8 which can be written 4(m-2) = 4m -8

OpenStudy (phi):

the idea is to rewrite 4m-8 as 4(m-2)

OpenStudy (stuck-help):

i cant understant what your doing i need you to show me in my problem i am very dilexix and cant put two problems together

OpenStudy (phi):

for the bottom (4 - 2m) factor out -2

OpenStudy (phi):

***i cant understant what your doing*** do you agree 4(m-2) = 4m-8 ?

OpenStudy (stuck-help):

yes so 4(2+m)=6+4m

OpenStudy (phi):

4(2+m) is 4*2 + 4*m or 8 +4m (not 6+4m)

OpenStudy (stuck-help):

sorry i added for some reason

OpenStudy (phi):

your problem has 4m-8 in the "top" we can rewrite the top as 4(m-2) now we look at the bottom (4-2m) factor out -2 (divide each term inside by -2) (4/-2 + -2/-2 m) and then multiply the whole thing by -2: -2(-2 + m) or -2(m-2)

OpenStudy (phi):

you problem becomes \[ \frac{4m-8}{4-2m} = \frac{4(m-2)}{-2(m-2)} \]

OpenStudy (phi):

if m is 2, the bottom becomes 0 (because (2-2) is zero). we do not allow divide by 0. so write down: \( m \ne 2 \) (don't allow m to be 2) if m-2 is not zero, we can simplify (m-2)/(m-2) to 1 (anything divided by itself is 1) so \[ \frac{4m-8}{4-2m} = \frac{4(m-2)}{-2(m-2)} = \frac{4}{-2} = -2 , m \ne 2\]

OpenStudy (stuck-help):

thank you